Iodine molecules dissociates into atoms after absorbing light of 4500 angstrom.If
1 quantum of radiation is absorbed by each molecule,calculate the kinetic energy
of iodine atoms.(bond energy of iodine molecules=240kJ/mol)

Grade:11

1 Answers

ABHISHEK JAIN AskiitiansExpert-IITD
13 years ago

Dear Shambhavi,

 Energy given to Iodine molecule = hc/λ

                                            = 6.62 × 10^-34 × 3 × 10^8 ÷ 4500× 10^-10

                                             =4.417 × 10^-19 J

also, energy used for braking up iodine molecule = 240 × 10^3 ÷ 6.023 × 10 ^23 =3.984× 10^-19 J

(bcoz B.E. is given in per mole and we have to find out per molecule)

 

Energy used in imperting Kinetic energy to two I atoms = [4.417 -3.984] × 10 ^-19 J

so K.E per iodine atom = [4.417-3.984/2] × 10^-19

                                 = .216 × 10^-19 J